DIFFERENTIAL GEOMETRY/PDE SEMINAR

Tuesday, July 29, 2003 Padelford C-36 1:30 pm (Note: Special Day and Time!)

Higher symmetries of the Laplacian

Michael G. Eastwood (University of Adelaide)

Which linear differential operators preserve harmonic functions? Even on Euclidean space, this is a deceptively simple question. The answer may be expressed in terms of conformal geometry and the AdS/CFT correspondence.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.